# DisasterM3: A Remote Sensing Vision-Language Dataset for Disaster Damage Assessment and Response

### v1

@@ 1 Introduction
Onset natural and man-made disasters represent one of humanity’s greatest challenges, causing devastating impacts across national borders [46, 8]. These catastrophic events (including earthquakes, tsunamis, floods, explosions, storms, etc) claim tens of thousands of lives globally each year while causing massive infrastructure damage and economic losses [33, 25]. Remote sensing (RS), as an ultra-long-distance Earth observation technology, has been widely used in disaster scenarios, i.e, hurricane damage assessment [29], landslide detection [36], mapping of burn area and ecological impacts [26], etc. Considering the urgency and timeliness of disaster relief, developing AI-based algorithms is necessary.
The recent advent of large vision-language models (VLMs) [16, 42, 6] has achieved substantial milestones in computer vision due to their exceptional ability to reason about visual and linguistic clues and summarize high-level human-readable text. Inspired by the success of the generic domain, remote sensing has also explored the applications of VLMs, i.e., image classification [14], image captioning [13], visual question answering [41], etc. These remote sensing-tailored VLMs show great potential as general-purpose task solvers for multi-task scenarios. Unlike existing research that primarily addresses general geospatial tasks, our work explores the reasoning capabilities of VLMs in extreme disaster scenarios, thereby supporting rescue teams and planning personnel in making informed decisions.
@@ 2 Related Work
General Vision-language Model. Assisted by the strong reasoning abilities of large language models, VLMs have transformed the visual perception domain by enabling the interpretation and reasoning about images through natural language interfaces. Several leading VLMs, including Flamingo [1], MiniGPT-4 [51], LLaVA [17], LLaVA-OneVision [16], InstructBLIP [6], and Qwen2-VL [42], have achieved remarkable results on vision-language tasks. However, these models are limited to generating only textual outputs that describe the image holistically. This restricts their applicability in damage assessment tasks that require the pixel-level detailed understanding. Several approaches have emerged to extend VLMs with fine-grained visual understanding. Ferret [47], Kosmos-2 [27], and VisionLLM [43] incorporate grounding functionalities through bounding box coordinate regression. Besides, LISA [15], PixelLM [31], GLaMM [30], and PerceptionGPT [28], integrate mask decoders to generate object masks from specialized tokens. For richer representation, PSALM [50] and HyperSeg [44] leverage queries in Mask2Former for unified segmentation. Despite their capabilities, generic VLMs exhibit substantial limitations in disaster scenarios due to insufficient domain-specific knowledge, restricting their operational utility in emergency response applications.
@@
Dataset Propose #Optical #SAR #MT pairs #Text Recognition Counting Localization Reasoning Caption
RSICD [23] General 10,921 - - 54,605 -
RSICap [11] General 2,585 - - 2,585 -
DIOR-RSVG [49] General 17,402 - - 38,320 Pixel
RRSIS-D [20] General 17,402 - - 17,402 Pixel
RSVQA-HR [22] General 10,659 - - 1,066,316 -
EarthVQA [41] General 6,000 - - 208,593 -
RSIEval [11] General 100 - - 933 -
VRSBench [18] General 29,614 - - 205,317 Box
XLRSBench [39] General 1,400 - - 45,942 Box
GeoChatSet [14] General 106,747 - - 308,861 Box
TeoChatlas [13] General 351,957 - 245,210 554,071 Box
FloodNet [29] Disaster 2,348 - - 7,345 -
@@
Remote Sensing Vision-language Dataset. Following the substantial progress of general VLMs, the RS field has likewise undergone accelerated development, accompanied by the emergence of numerous specialized vision-language datasets. Focusing on holistic analysis, EarthVQA [41] and RSIEval [11] datasets provide manual instructions for visual question answering (VQA) and image captioning tasks. Leveraging GPT-4, VRSBench [18] introduced visual grounding tasks to evaluate the object reasoning abilities and XLRSBench [39] focuses on ultra-high-resolution image understanding. GeoChatSet [14] and TeoChatlas [13] collect the existing classification and detection datasets for secondary development, formulating the unified instruction-following datasets. Although TeoChatlas involves some disaster scenes, the instructions focus on common recognition tasks. FloodNet [29] is a VQA disaster dataset that assesses the buildings and roads affected by Hurricane Harvey. Limited by its single disaster and simple tasks, it is difficult to fully unleash the potential of VLMs. Overall, RS visual-language datasets for general geospatial tasks have reached a considerable level of maturity, yet there persists a notable deficiency in datasets addressing specialized geoscience challenges. For this case, we design the DisasterM3 dataset that is tailored for global disaster assessment and response with multi-sensor images, bi-temporal inputs, refined damage masks, and diverse visual understanding tasks in the context of disaster.
@@ 3 DisasterM3 Dataset
As shown in Fig. 2, we collect 36 historical natural and man-made significant disasters to construct the DisasterM3 dataset. There are 26 events from the xBD [9] and BRIGHT [5] dataset, we extend 10 new events using Maxar’s Open Data program [24]. Considering these optical sensors (WorldView series) have similar spatial resolutions, all pre- and post-disaster images were pre-processed into 0.8m. We collect the post-disaster Synthetic Aperture Radar (SAR) images from Capella Space [4] and Umbra [37]. Considering the amplitude data in the VV or HH bands, SAR images were terrain-corrected, stretched into [0, 255], and finally resampled to match the optical resolution. We performed the georeferencing to ensure that the pre- and post-disaster image pairs are strictly aligned spatially. Following the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products [38], and the Federal Emergency Management Agency (FEMA) [7], we design 9 essential tasks required for disaster assessment and response, evaluating the VLM performances from different aspects.
@@ 3.1 Perception and Reasoning Tasks in the Context of Disaster
Disaster Recognition. The disaster recognition tasks provide a brief description of disaster scenes, i.e., disaster types, land-use types and key disaster-bearing bodies. The disaster type follows the official definition and and we chose 13 common land-use types from the AID dataset [45] for annotation. The land-use answers include: airport, bridge, river, forest, low vegetation, pond, parking, port, viaduct, residential area, industrial area, commercial area, and sea. Disaster-bearing bodies are the key resources that are damaged by disasters [8], and we focus on 12 types, i.e., building, stadium, open-space ground, bridge, dam, road, port facility, storage tank, farmland, forest, coastline, and mining area. Based on basic recognition types, users could have a rough disaster profile.
Damage Assessment. The damage assessment provides a quantitative analysis of disaster-bearing bodies. We chose the road and building, two important man-made structures for damage assessment. We annotate instance-level building damage masks using ‘intact’, ‘damaged’, and ‘destroyed’ types following FEMA guidelines. As a critical transportation hub, road accessibility plays a vital role in emergency response and recovery efforts. We classify the damaged roads into three types, i.e., ‘intact’, ‘flooded (blocked by water)’, and ‘debris covered (blocked by debris)’. Based on these damage masks, the building counting and road area estimation instructions were automatically generated. The imbalanced sample distributions of damaged buildings and roads (Fig. 2) reveal the actual challenges for model optimization.
@@
Disaster Referring Segmentation. Each disaster includes different forming factors and prone environments. In addition to disaster-bearing-body mapping, we identify the key visual objects and perform risk analysis using referring segmentation. As shown in Fig. 3, the first example shows an earthquake scene. In addition to referring segmentation for disaster-bearing bodies, we also design the task for finding the optimal rescue places shown in Fig 3(d). Similarly, Fig 3(e) shows the place where rescuers could find the available vehicles for dispatch. As for the volcano eruption scene, we set the instruction tasks to individually map damaged buildings and roads, as well as the lava. Considering the situation, the intact buildings near the lava are also required for segmentation. By polygon distance analysis using the ArcGIS toolbox, the intact buildings within a 100-meter proximity to lava are segmented, providing early warning information. All the referring segmentation tasks are designed according to the specific disaster scenarios, which enable the rescuers to accurately locate the disaster-related objects and places.
@@
Disaster Comprehensive Report. To go beyond traditional perception tasks, the comprehensive reports are designed for the holistic analysis of disaster situations. Fig. 5 shows two samples of disaster caption and recovery advice. The earthquake caption describes the collapsed buildings and blocked roads, causing severe traffic congestion. Immediate response advice prioritizes the deployment of temporary shelters within the stadium for displaced survivors, a recommendation visibly implemented in the post-disaster image. Long-term recovery focuses on earthquake-resistant strategies in rebuilding and disaster protocols to mitigate seismic risks. The flooding caption summarizes that roads, buildings, and natural areas experienced severe inundation, while water bodies expanded and merged with flooded regions. Correspondingly, repairing critical transportation infrastructure, establishing temporary residential facilities, and implementing disease prevention protocols are proposed as immediate response measures. The installation of drainage systems integrated with local hydrological networks is recommended as a long-term strategy. Fig. 5 (e) and (f) shows the word cloud of disaster reports. Thanks to the wide range of disaster types, the words are diverse in terms of both nouns and verbs. Most words are disaster-centric, describing bearing bodies, damage impacts, response strategies, etc. Comprehensive disaster reports equip rescuers with enhanced situational awareness and evidence-based decision support.
@@ 3.2 Dataset Construction Pipeline
Following the common vision-language data pipeline [21, 48], we divided the whole dataset into Instruct (17,190 Optical images, 3,798 SAR images, and 92,968 instruction pairs) and Bench sets (5,024 Optical images, 976 SAR images, and 30,042 instruction pairs). We describe the detailed annotation process in Fig. 6. As for recognition tasks, GPT-4o was employed to generate diverse prompt variations with similar semantic intent. Disaster domain experts subsequently annotate correct answers for these prompts. These question-answer pairs constitute the Instruct training set. To formulate the multiple-choice Bench set, correct answers were combined with other options. Regarding counting tasks, we counted semantic polygons using annotated road and building damage masks, generating correct answers. The similar options are generated with controlled deviations (±20% and ±40%) to maintain plausibility. As for relational reasoning, the experts annotate bounding boxes and describe the concrete relationship. We use GPT4-o to analyze the image by listing other significant relationships, generating alternatives. As for disaster reports, by referring to bi-temporal images and all basic task information, multiple experts draft the disaster caption and restoration advice following goals of UNITAR and and FEMA projects. GPT4 then polished the reports and corrected grammar errors. Finally, the multi-round verification was performed for controlling quality (Appendix § A). As for uninterpretable parts of SAR imagery, we annotate answers using co-registered optical images and then apply the instructions to SAR images. Using this pipeline, more future disasters can be effectively extended for DisasterM3 dataset.
@@ 4 Benchmark Experiments
Implementation Setting. As the DisasterM3 dataset features multi-sensor and multi-task, we comprehensively benchmark VLMs under four settings: Optical-Optical and Optical-SAR QA tasks, as well as Optical-Optical and Optical-SAR referring segmentation tasks.. As for QA tasks, LLaVA-OneVision [16], InternVL3 [53], Kimi-VL [35], and Qwen2.5-VL [3]. In addition, we also tested commercial models such as GPT4o [12] and Claude-3 [2] for comparison with the open-source models. As for remote sensing VLMs GeoChat [14], TeoChat [13], EarthDial [34] are chosen for evaluation. As for referring segmentation models, generic VLMs models such as LISA [15], PSALM [50], and HyperSeg [44], alongside the remote sensing model GeoPixel [32] were benchmarked. We fine-tuned Qwen2.5-VL-7B, InvernVL3-8B, LISA and PSALM on our Instruct set. Model details are provided in Appendix § B.
Evaluation Metrics. Following common settings [16, 34], we adopted accuracy (%) for the multiple-choice tasks, i.e., disaster scene recognition (DSR), disaster type recognition (DTR), bearing body recognition (BBR), damaged building counting (DBC), damaged road estimation (DRE), object relational reasoning (ORR). The open-ended tasks are scored using GPT4.1 at a scale of 5 points. Disaster caption is measured from damage assessment precision (DAP), damage detail recall (DDR), and factual correctness (FC). Restoration advice is measured from recovery necessity (RN), strategic completeness (SC), and action priority precision (APP). The average accuracy (AVG) denotes the overall performance. Evaluation prompts are provided in Appendix § C. As for referring segmentation, we chose cIoU and mIoU following previous work [15, 50].

### v2

# Hongruixuan Chen1, Jian Song2, Junshi Xia2, Zhuo Zheng5, Naoto Yokoya1,2

@@ 1 Introduction
Onset natural and man-made disasters represent one of humanity’s greatest challenges, causing devastating impacts across national borders xu2025implementing ; frankenberg2020effects . These catastrophic events (including earthquakes, tsunamis, floods, explosions, storms, etc) claim tens of thousands of lives globally each year while causing massive infrastructure damage and economic losses shugar2021massive ; palmer2020new . Remote sensing (RS), as an ultra-long-distance Earth observation technology, has been widely used in disaster scenarios, i.e., hurricane damage assessment rahnemoonfar2021floodnet , landslide detection tingzon2023fusing , mapping of burn area and ecological impacts pan2022deep , etc. Considering the urgency and timeliness of disaster relief, developing AI-based algorithms is necessary.
The recent advent of large vision-language models (VLMs) li2024llava-ov ; wang2024qwen2 ; dai2023instructblip has achieved substantial milestones in computer vision due to their exceptional ability to reason about visual and linguistic clues and summarize high-level human-readable text. Inspired by the success of the generic domain, remote sensing has also explored the applications of VLMs, i.e., image classification kuckreja2024geochat , image captioning irvin2024teochat , visual question answering wang2024earthvqa , etc. These remote sensing-tailored VLMs show great potential as general-purpose task solvers for multi-task scenarios. Unlike existing research that primarily addresses general geospatial tasks, our work explores the reasoning capabilities of VLMs in extreme disaster scenarios, thereby supporting rescue teams and planning personnel in making informed decisions.
@@ 2 Related Work
General Vision-language Model. Assisted by the strong reasoning abilities of large language models, VLMs have transformed the visual perception domain by enabling the interpretation and reasoning about images through natural language interfaces. Several leading VLMs, including Flamingo alayrac2022flamingo , MiniGPT-4 zhuminigpt , LLaVA li2024llava , LLaVA-OneVision li2024llava-ov , InstructBLIP dai2023instructblip , and Qwen2-VL wang2024qwen2 , have achieved remarkable results on vision-language tasks. However, these models are limited to generating only textual outputs that describe the image holistically. This restricts their applicability in damage assessment tasks that require the pixel-level detailed understanding. Several approaches have emerged to extend VLMs with fine-grained visual understanding. Ferret youferret , Kosmos-2 peng2023kosmos , and VisionLLM wang2023visionllm incorporate grounding functionalities through bounding box coordinate regression. Besides, LISA lai2024lisa , PixelLM ren2024pixellm , GLaMM rasheed2024glamm , and PerceptionGPT pi2024perceptiongpt , integrate mask decoders to generate object masks from specialized tokens. For richer representation, PSALM zhang2025psalm and HyperSeg wei2024hyperseg leverage queries in Mask2Former for unified segmentation. Despite their capabilities, generic VLMs exhibit substantial limitations in disaster scenarios due to insufficient domain-specific knowledge, restricting their operational utility in emergency response applications.
@@
Dataset Propose #Optical #SAR #MT pairs* #Text Recognition Counting Localization Reasoning Caption
RSICD lu2017exploring General 10,921 - - 54,605 -
RSICap hu2023rsgpt General 2,585 - - 2,585 -
DIOR-RSVG zhan2023rsvg General 17,402 - - 38,320 Box
RRSIS-D liu2024rotated General 17,402 - - 17,402 Pixel
RSVQA-HR lobry2020rsvqa General 10,659 - - 1,066,316 -
EarthVQA wang2024earthvqa General 6,000 - - 208,593 -
RSIEval hu2023rsgpt General 100 - - 933 -
VRSBench li2024vrsbench General 29,614 - - 205,317 Box
XLRSBench wang2025xlrsbench General 1,400 - - 45,942 Box
GeoChatSet kuckreja2024geochat General 106,747 - - 308,861 Box
TeoChatlas irvin2024teochat General 351,957 - 245,210 554,071 Box
FloodNet rahnemoonfar2021floodnet Disaster 2,348 - - 7,345 -
@@
    MT pairs (multi-temporal pairs) denote the number of pre/post-disaster image pairs.
Remote Sensing Vision-language Dataset. Following the substantial progress of general VLMs, the RS field has likewise undergone accelerated development, accompanied by the emergence of numerous specialized vision-language datasets. Focusing on holistic analysis, EarthVQA wang2024earthvqa and RSIEval hu2023rsgpt datasets provide manual instructions for visual question answering (VQA) and image captioning tasks. Leveraging GPT-4, VRSBench li2024vrsbench introduced visual grounding tasks to evaluate the object reasoning abilities and XLRSBench wang2025xlrsbench focuses on ultra-high-resolution image understanding. GeoChatSet kuckreja2024geochat and TeoChatlas irvin2024teochat collect the existing classification and detection datasets for secondary development, formulating the unified instruction-following datasets. Although TeoChatlas involves some disaster scenes, the instructions focus on common recognition tasks. FloodNet rahnemoonfar2021floodnet is a VQA disaster dataset that assesses the buildings and roads affected by Hurricane Harvey. Limited by its single disaster and simple tasks, it is difficult to fully unleash the potential of VLMs. Overall, RS visual-language datasets for general geospatial tasks have reached a considerable level of maturity, yet there persists a notable deficiency in datasets addressing specialized geoscience challenges. For this case, we design the DisasterM3 dataset that is tailored for global disaster assessment and response with multi-sensor images, bi-temporal inputs, refined damage masks, and diverse visual understanding tasks in the context of disaster.
@@ 3 DisasterM3 Dataset
As shown in Fig. 2, we collect 36 historical natural and man-made significant disasters to construct the DisasterM3 dataset. There are 26 events from the xBD gupta2019xbd and BRIGHT chen2025bright dataset, we extend 10 new events using Maxar’s Open Data program maxar2025 . Considering these optical sensors (WorldView series) have similar spatial resolutions, all pre- and post-disaster images were pre-processed into 0.8 m. We collect the post-disaster Synthetic Aperture Radar (SAR) images from Capella Space capella2025 and Umbra umbra2025 . Considering the amplitude data in the VV or HH bands, SAR images were terrain-corrected, stretched into [0, 255], and finally resampled to match the optical resolution. We performed the georeferencing to ensure that the pre- and post-disaster image pairs are strictly aligned spatially. Following the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products unosat , and the Federal Emergency Management Agency (FEMA) emergency2002federal , we design 9 essential tasks required for disaster assessment and response, evaluating the VLM performances from different aspects.
@@ 3.1 Perception and Reasoning Tasks in the Context of Disaster
Disaster Recognition. The disaster recognition tasks provide a brief description of disaster scenes, i.e., disaster types, land-use types and key disaster-bearing body. The disaster type follows the official definition and we chose 13 common land-use types from the AID dataset xia2017aid for annotation. The land-use answers include: airport, bridge, river, forest, low vegetation, pond, parking, port, viaduct, residential area, industrial area, commercial area, and sea. Disaster-bearing bodies are the key resources that are damaged by disasters frankenberg2020effects , and we focus on 12 types, i.e., building, stadium, open-space ground, bridge, dam, road, port facility, storage tank, farmland, forest, coastline, and mining area. Based on basic recognition types, users could have a rough disaster profile.
Damage Assessment. The damage assessment provides a quantitative analysis of disaster-bearing body. We chose the road and building, two important man-made structures for damage assessment. We annotate instance-level building damage masks using ‘intact’, ‘damaged’, and ‘destroyed’ types following FEMA guidelines. As a critical transportation hub, road accessibility plays a vital role in emergency response and recovery efforts. We classify the damaged roads into three types, i.e., ‘intact’, ‘flooded (blocked by water)’, and ‘debris covered (blocked by debris)’. Based on these damage masks, the building counting and road area estimation instructions were automatically generated. The imbalanced sample distributions of damaged buildings and roads (Fig. 2) reveal the actual challenges for model optimization.
@@
Disaster Referring Segmentation. Each disaster includes different forming factors and prone environments. In addition to disaster-bearing-body mapping, we identify the key visual objects and perform risk analysis using referring segmentation. As shown in Fig. 3, the first example shows an earthquake scene. In addition to referring segmentation for disaster-bearing body, we also design the task for finding the optimal rescue places shown in Fig 3(d). Similarly, Fig 3(e) shows the place where rescuers could find the available vehicles for dispatch. As for the volcano eruption scene, we set the instruction tasks to individually map damaged buildings and roads, as well as the lava. Considering the situation, the intact buildings near the lava are also required for segmentation. By polygon distance analysis using the ArcGIS toolbox, the intact buildings within a 100-meter proximity to lava are segmented, providing early warning information. All the referring segmentation tasks are designed according to the specific disaster scenarios, which enable the rescuers to accurately locate the disaster-related objects and places.
@@
Disaster Comprehensive Report. To go beyond traditional perception tasks, the comprehensive reports are designed for the holistic analysis of disaster situations. Fig. 5 shows two samples of disaster caption and restoration advice. The earthquake caption describes the collapsed buildings and blocked roads, causing severe traffic congestion. Immediate response advice prioritizes the deployment of temporary shelters within the stadium for displaced survivors, a recommendation visibly implemented in the post-disaster image. Long-term recovery focuses on earthquake-resistant strategies in rebuilding and disaster protocols to mitigate seismic risks. The flooding caption summarizes that roads, buildings, and natural areas experienced severe inundation, while water bodies expanded and merged with flooded regions. Correspondingly, repairing critical transportation infrastructure, establishing temporary residential facilities, and implementing disease prevention protocols are proposed as immediate response measures. The installation of drainage systems integrated with local hydrological networks is recommended as a long-term strategy. Fig. 5 (e) and (f) shows the word cloud of disaster reports. Thanks to the wide range of disaster types, the words are diverse in terms of both nouns and verbs. Most words are disaster-centric, describing bearing bodies, damage impacts, response strategies, etc. Comprehensive disaster reports equip rescuers with enhanced situational awareness and evidence-based decision support.
@@ 3.2 Dataset Construction Pipeline
Following the common vision-language data pipeline liu2024bench ; Yue_2024_CVPR , we divided the whole dataset into Instruct (17,190 Optical images, 3,798 SAR images, and 92,968 instruction pairs) and Bench sets (5,024 Optical images, 976 SAR images, and 30,042 instruction pairs). We describe the detailed annotation process in Fig. 6. As for recognition tasks, GPT-4o was employed to generate diverse prompt variations with similar semantic intent. Disaster domain experts subsequently annotate correct answers for these prompts. These question-answer pairs constitute the Instruct training set. To formulate the multiple-choice Bench set, correct answers were combined with other options. Regarding counting tasks, we counted semantic polygons using annotated road and building damage masks, generating correct answers. The similar options are generated with controlled deviations (±20% and ±40%) to maintain plausibility. As for relational reasoning, the experts annotate bounding boxes and describe the concrete relationship. We use GPT-4o to analyze the image by listing other significant relationships, generating alternatives. As for disaster reports, by referring to bi-temporal images and all basic task information, multiple experts draft the disaster caption and restoration advice following goals of UNITAR and FEMA projects. GPT-4o then polished the reports and corrected grammar errors. Finally, the multi-round verification was performed for controlling quality (Appendix §A). As for uninterpretable parts of SAR imagery, we annotate answers using co-registered optical images and then apply the instructions to SAR images. Using this pipeline, more future disasters can be effectively extended for DisasterM3 dataset.
@@ 4 Benchmark Experiments
Implementation Setting. As the DisasterM3 dataset features multi-sensor and multi-task, we comprehensively benchmark VLMs under four settings: Optical-Optical and Optical-SAR QA tasks, as well as Optical-Optical and Optical-SAR referring segmentation tasks.. As for QA tasks, LLaVA-OneVision li2024llava-ov , InternVL3 zhu2025internvl3 , Kimi-VL team2025kimi , and Qwen2.5-VL bai2025qwen2 . In addition, we also tested commercial models such as GPT-4o hurst2024gpt and Claude-3 anthropic2024claude3 for comparison with the open-source models. As for remote sensing VLMs GeoChat kuckreja2024geochat , TeoChat irvin2024teochat , EarthDial soni2024earthdial are chosen for evaluation. As for referring segmentation models, generic VLMs models such as LISA lai2024lisa , PSALM zhang2025psalm , and HyperSeg wei2024hyperseg , alongside the remote sensing model GeoPixel shabbir2025geopixel were benchmarked. We fine-tuned Qwen2.5-VL-7B, InternVL3-8B, LISA and PSALM on our Instruct set. Model details are provided in Appendix §B.
Evaluation Metrics. Following common settings li2024llava-ov ; soni2024earthdial , we adopted accuracy (%) for the multiple-choice tasks, i.e., disaster scene recognition (DSR), disaster type recognition (DTR), bearing body recognition (BBR), damaged building counting (DBC), damaged road estimation (DRE), object relational reasoning (ORR). The open-ended tasks are scored using GPT-4.1 at a scale of 5 points. Disaster caption is measured from damage assessment precision (DAP), damage detail recall (DDR), and factual correctness (FC). Restoration advice is measured from recovery necessity (RN), strategic completeness (SC), and action priority precision (APP). The average accuracy (AVG) denotes the overall performance. Evaluation prompts are provided in Appendix §C. As for referring segmentation, we chose cIoU and mIoU following previous work lai2024lisa ; zhang2025psalm .